A force acts for 10s on a body of man 10kg initially at rest ,after which the force ceases to act. if the body traversed 0.5m in the next 5s then find F.

Grade:11

1 Answers

DR STRANGE
7 years ago
HI Sai,
let the force be F,
F=Ma
a=F/M=F/10
NOW, v=u+at, v= 0+(F/10)*10=F (u=0, initially at rest)
now, v=s/t.
it traversed 0.5m in next 5s
v=F=0.5/5=0.1N............ANS
Plz approve if you are satisfied :)